OpenStudy (firejay5):

While at Magic Springs, Katie goes down a frictionless water slide that is 9.0 m above the ground. At the bottom of the slide, Katie reaches a final speed of 13 m/s. What was Katie's initial speed when she pushed off the top of the slide?

OpenStudy (firejay5):

@ganeshie8

OpenStudy (anonymous):

Consider this problem in terms of conservation of energy\[KE _{Top}+PE _{Top}=KE _{Bottom}+PE _{Bottom}\]where KE is the kinetic energy and PE is the potential energy.

OpenStudy (anonymous):

i believe he has the formula wrong, its stupid - girlfriend = nerd

OpenStudy (anonymous):

jk jk sorrie

OpenStudy (perl):

do you need help with Kinetic energy. KE = 1/2 mv^2 PE = m*g*h